### v2

# Boosting Memory Efficiency in Transfer Learning for High-Resolution Medical Image Classification

Abstract
The success of large-scale pre-trained models has established fine-tuning as a standard method for achieving significant improvements in downstream tasks. However, fine-tuning the entire parameter set of a pre-trained model is costly. Parameter-efficient transfer learning (PETL) has recently emerged as a cost-effective alternative for adapting pre-trained models to downstream tasks. Despite its advantages, the increasing model size and input resolution present challenges for PETL, as the training memory consumption is not reduced as effectively as the parameter usage. In this paper, we introduce Fine-grained Prompt Tuning plus (FPT+), a PETL method designed for high-resolution medical image classification, which significantly reduces the training memory consumption compared to other PETL methods. FPT+ performs transfer learning by training a lightweight side network and accessing pre-trained knowledge from a large pre-trained model (LPM) through fine-grained prompts and fusion modules. Specifically, we freeze the LPM of interest and construct a learnable lightweight side network. The frozen LPM processes high-resolution images to extract fine-grained features, while the side network employs corresponding down-sampled low-resolution images to minimize the memory usage. To enable the side network to leverage pre-trained knowledge, we propose fine-grained prompts and fusion modules, which collaborate to summarize information through the LPM’s intermediate activations. We evaluate FPT+ on eight medical image datasets of varying sizes, modalities, and complexities. Experimental results demonstrate that FPT+ outperforms other PETL methods, using only 1.03% of the learnable parameters and 3.18% of the memory required for fine-tuning an entire ViT-B model. Our code is available online.
Index Terms:
Parameter-efficient transfer learning, Memory-effificent transfer learning, Large-scale pre-trained models, High-resolution medical image classification.
I Introduction
Recently, large pre-trained models (LPMs) have made remarkable achievements across diverse domains [38, 13, 49, 37, 52, 29, 8]. Utilizing the technique of transfer learning [47], pre-trained models can be effectively adapted to specific downstream tasks by initializing task-specific models with weights from a pre-trained model of interest, followed by training on task-specific datasets. However, as model sizes grow rapidly (e.g. ViT-B, a variant of vision transformer (ViT) [15], has 86 million parameters), fine-tuning the entire parameter set of an LPM has become very costly, demanding substantial training resources such as GPU memory and significantly limiting its practical applicability.
Figure 1: We utilize vision transformer (ViT-B) [15], pre-trained on ImageNet-21K [40], as the encoder. The reported performance is the average of performance metrics across eight medical image datasets. The size of the dot denotes the memory consumption during transfer learning. Our proposed method, FPT+, distinguishes itself from other PETL methods by achieving the best balance between performance and efficiency.
In response, the concept of PETL [14] has emerged. PETL offers a strategic approach for transferring pre-trained models by selectively updating a small subset of the pre-trained parameters or introducing a modest number of additional parameters specific to the new task while freezing the majority of the pre-trained parameters.
PETL initially gained substantial attention within the natural language processing (NLP) community before finding its application in computer vision (CV) where various PETL methods have shown effectiveness across diverse tasks and datasets. While these approaches have achieved notable success, they were primarily designed for natural images and the field of medical image analysis has not yet fully benefited from such advances [16]. The unique challenges of medical image analysis require specialized PETL methods due to the domain gap between natural and medical images [41, 57, 28].
In natural images, the objects of interest typically occupy a significant portion of the image and exhibit clear and noticeable attributes (e.g., shape and color). In contrast, the diagnostic cues within medical images are often tiny and spread throughout the entire visual field, making fine-grained information crucial for accurate medical image analysis [25]. Providing such fine-grained information often requires high-resolution input images [24]. However, as illustrated in Fig. 2, this preference for high resolution comes at the cost of increased GPU memory consumption and higher training expenses. For example, a ViT with a patch size of 16 divides a 224 $\times$ 224 image into 196 patches, whereas a 512 $\times$ 512 image generates 1024 patches. This sharp rise in computational demand makes fine-tuning LPMs both expensive and, in some cases, infeasible on devices with limited GPU memory.
Moreover, it is worth noting that existing LPMs are primarily pre-trained on large natural image datasets like ImageNet [12]. Attempting to bridge the knowledge gap between natural and medical images by merely updating a small subset of the pre-trained parameters or introducing additional parameters within a frozen LPM may not be sufficient to address the significant differences between these two domains [57].
In this work, we propose a novel parameter and memory-efficient PETL method, namely Fine-grained Prompt Tuning plus (FPT+). FPT+ aims to enhance the efficiency of PETL specifically for high-resolution medical images. Existing PETL methods typically involve training a subset of parameters within the LPM of interest. However, updating these parameters still requires gradients’ computation through the frozen layers during back-propagation, significantly increasing memory consumption. To address this issue, FPT+ employs a lightweight additive network inspired by the concept of a side network [50, 44]. This design treats the entire pre-trained model as a frozen feature extractor, eliminating the need for back-propagation through the LPM. To leverage pre-trained knowledge from the LPM, FPT+ introduces the concept of fine-grained prompts and fine-grained fusion modules as the bridging components. Fine-grained prompts consist of a small set of trainable embeddings prepended to the input sequence of the side network. Instead of engaging the entire input sequence for interaction with the fine-grained features from the pre-trained model, the fusion process involves only the fine-grained prompts through cross-attention mechanisms. These prompts subsequently convey fine-grained information to the side network during forward propagation, allowing the integration of pre-trained knowledge into the learnable side network at a low cost.
Figure 2: High resolution comes at the cost of heightened GPU memory consumption. The red line represents the maximum memory size of the NVIDIA RTX A6000 GPU used in our experiments. Our proposed FPT+ makes training on super high-resolution images feasible.
Additionally, considering that long input sequences from high-resolution images result in substantial intermediate activations, which can be computationally expensive, FPT+ employs important tokens selection to mitigate this issue. Based on the prior knowledge that medical images of the same modality often share similar anatomical characteristics and have small objects of interest (e.g., lesions and vessels in fundus images), FPT+ selectively fuses features from the intermediate activations to reduce GPU memory consumption. Preloading techniques are also proposed to improve the training efficiency by pre-storing the LPM’s intermediate features.
To validate the effectiveness of the proposed FPT+, we adopt eight publicly accessible medical image datasets with varying sizes, modalities, and complexities. Experimental results demonstrate that FPT+ achieves the best performance-efficiency balance, outperforming other state-of-the-art (SOTA) PETL methods with only 1.03% of the trainable parameters and 3.18% of the memory cost required for fine-tuning an entire model.
This paper’s main contributions are summarized as follows:
  • We propose a novel PETL method, namely FPT+, for high-resolution medical image classification. FPT+ maximizes the memory efficiency of transfer learning through asymmetric input and important tokens selection. The effectiveness and efficiency of adapting out-of-domain knowledge are achieved by our fine-grained prompts and fusion modules.
  • FPT+ sets a new benchmark in memory efficiency for PETL methods, achieving competitive performance utilizing only 3.18% of the memory cost required for fine-tuning an entire model. As shown in Fig. 2, FPT+ significantly extends the applicability of transfer learning within the realm of high-resolution inputs.
  • Extensive experiments are conducted on eight medical image datasets with various sizes, modalities and complexities, demonstrating the effectiveness of FPT+ by achieving the best balance between performance and parameter/memory efficiency. Additionally, resource consumption is comprehensively evaluated to establish the superior training efficiency of FPT+.
  • We evaluate different LPMs for transfer learning, including both supervised and self-supervised methods, emphasizing the generalizability of FPT+ across different LPM scales and different pre-trained weights.
This paper extends our previous work FPT [23]. We systematically investigate the impact of several key components in FPT and refine the entire framework to make it simpler and more efficient. Specifically, our extension is presented in several aspects: (1) We reduce the number of layers in the side network and its corresponding fusion modules to significantly decrease memory consumption; (2) We redesign the fusion module by replacing in-projector with layer normalization and align the initial hidden dimension of the fine-grained prompts with the frozen LPM; (3) All fusion modules share the same weights for the out-projector, further reducing the number of learnable parameters; (4) The new design improves the effectiveness of learning pre-trained knowledge, allowing smaller low-resolution inputs for the side network, thus further reducing memory consumption; (5) To verify the generalizability of FPT+, we validate the framework using four additional datasets; (6) We evaluate the scalability and feasibility of FPT+ across pre-trained models of different scales and different weights; (7) Additional efficiency metrics, including training time, throughput, and power usage, are evaluated to comprehensively assess FPT+’s resource consumption and training efficiency; (8) Extensive ablation studies are conducted to qualitatively and quantitatively analyze the impact of each component in FPT+ on performance, parameter usage and memory efficiency.
II Related Work
II-A Parameter-efficient Transfer Learning
In recent years, the practice of fine-tuning expansive pre-trained models has shown encouraging potential across a range of tasks in the field of NLP. With the rapidly increasing sizes of pre-trained models, typically transformers [46, 3], the concept of PETL has emerged as a strategic solution to mitigate the computational burden in adapting large-scale pre-trained models to downstream tasks.
LoRA-based methods [22, 51] freeze pre-trained weights and inject trainable rank decomposition matrices into each layer of a transformer to reduce the number of trainable parameters. Prompt tuning [31] appends the input with a trainable tensor, termed prompt, while freezing all other parameters. Adapter [21] introduces the idea of adding new learnable modules among layers of a frozen pre-trained model to adapt to new downstream tasks. BitFit [2] tunes only the bias term of a pre-trained model to effectively adapt the model to downstream tasks.
Drawing inspiration from NLP advancements, large-scale pre-trained models for CV are also receiving growing attention. ViTs are widely adopted as the backbone due to their modeling power and scalability to model and data size [51, 56, 35, 4]. While most PETL methods were initially designed for transformers, recent works have demonstrated their effectiveness in ViTs as well.
There are also several PETL methods specifically designed for ViT [7, 27, 55, 34, 17]. Attention tuning [45] fine-tunes only the attention layers to adapt ViTs to various downstream tasks. Vision prompt tuning [27] expands upon the concept of prompt tuning by introducing learnable prompts into each block of a frozen pre-trained ViT. Side tuning [50] employs an additive side network, which combines features from a frozen pre-trained model using a simple summation operation facilitated by a trainable gate. Liu et al. [34] propose Explicit Visual Prompting to improve fine-tuning on segmentation tasks by taking the high-frequency components in the input as the prompts. CoOp [55] and CoCoOp [54] are proposed to specifically adapt CLIP-like vision-language pre-trained models to downstream image recognition tasks. Zhao et al. [53] leverage task-specific information to select salient channels and then fine-tune only those channels to enhance parameter efficiency.
Unlike other PETL methods, FPT [23] and our proposed FPT+ aim to improve both parameter and memory efficiency during transfer learning, which is a specific and common requirement in the field of medical image analysis.
II-B PETL for Medical Image Analysis
While several PETL methods have emerged in the realm of CV, the application of PETL to ViTs for medical images remains relatively limited, despite the efficiency requirement inherent in the medical domain [16]. Raman et al. [16] comprehensively evaluate the performance of different PETL methods for medical image analysis. Rodríguez et al. [43] study adapters designed for dense prediction tasks to improve medical image segmentation. Dynamic Visual Prompt Tuning [19] is designed to learn sample-specific and domain-specific features to enhance the fine-tuning performance in the medical domain. Zu et al. [57] propose Embedded Prompt Tuning by embedding prompt tokens into expanded channels for few-shot medical image classification tasks. Wu et al. [48] combine the SAM model [29] with adapters, which outperforms existing SOTA methods on several downstream medical image segmentation tasks.
To the best of our knowledge, our previous work, FPT [23], is the first PETL method specifically designed for high-resolution input settings, simultaneously enhancing both parameter and memory efficiency. We extend FPT by simplifying the entire framework to make it more efficient and more effective. Specifically, we halve the number of layers in the side network and the corresponding fusion modules, and redesign the fusion module’s in-projector. Unlike other PETL methods for medical image analysis, our proposed FPT+ framework aims to simultaneously address the challenge of unaffordable parameter as well as memory utilization when transferring LPMs to high-resolution medical image classification tasks.
Figure 3: Our FPT+ framework comprises the following components: (a) The large-scale pre-trained model remains frozen and receives high-resolution images as the input. A lightweight, learnable side network is employed to integrate knowledge from the pre-trained model into new downstream tasks through the utilization of fine-grained prompts and fusion modules. (b) The fine-grained fusion module operates by taking fine-grained prompts and important features extracted from the pre-trained model as inputs. Prompts and features are fused through feature alignment and cross-attention.
III Method
In this work, we propose FPT+, a parameter and memory efficient transfer learning method for high-resolution medical image classification. FPT+ can: 1) efficiently integrate fine-grained information from a pre-trained model into a new downstream model; and 2) effectively adapt knowledge from an out-of-domain pre-trained model. As shown in Fig. 3, a separate and lightweight side network is first employed to learn and infer at an affordable training cost in terms of both the number of trainable parameters and memory consumption. Then, fine-grained prompts and fusion modules are proposed to convey fine-grained pre-trained knowledge to the side network effectively and efficiently. Asymmetric input and important tokens selection are introduced to further reduce the computational burden. The motivation, benefits, and functioning of each component are delineated in subsequent sections.
III-A Vision Transformer
In this section, we provide a brief overview of ViT [15]. A conventional ViT architecture consists of a linear projection layer followed by multiple transformer blocks. For a given image $x\in\mathbb{R}^{H\times W\times C}$ with spatial dimensions $(H,W)$ and channel dimension $C$, the initial step involves partitioning the image into non-overlapping 2D patches of size $P\times P$. Here, $P$ represents the patch size specific to the ViT of interest. Subsequently, these 2D patches are flattened and transformed into a latent embedding space with hidden dimension $D$, facilitated by a linear projection layer. This process yields an input sequence $\bm{z}_{0}={z_{p}^{i}\in\mathbb{R}^{D}|i=1,...,N}$ with a length of $N=HW/P^{2}$. Each element in this input sequence is termed a token. The input is then processed by multiple transformer blocks, each consisting of a multi-headed self-attention (MSA) layer and a multi-layer perceptron (MLP) block:
$\bm{z}_{l}^{\prime}=\text{MSA}(\bm{z}_{l-1})+\bm{z}_{l-1},$ (1)
$\bm{z}_{l}=\text{MLP}(\bm{z}_{l}^{\prime})+\bm{z}_{l}^{\prime},$ (2)
where $l$ indexes layers of the ViT. The class token and layer normalization operations are omitted in the above formulas for simplification.
Within each self-attention layer, individual tokens in the input sequence are mapped to Q (query), K (key), and V (value) representations for computations in the attention module (Attn) [46] as follows:
$\text{MSA}(\bm{z})$ $=\text{Attn}(Q,K,V)$
$=\text{Attn}(f_{Q}(\bm{z}),f_{K}(\bm{z}),f_{V}(\bm{z})),$ (3)
where $f_{Q}$, $f_{K}$, and $f_{V}$ denote the mapping functions in the attention module.
III-B Side Tuning
Unlike popular PETL methods which introduce supplementary modules into transformer blocks, our approach, drawing inspiration from [50, 44], adopts a lightweight and separate side network. The side network takes advantages of the LPM by utilizing intermediate features from the frozen pre-trained model. As depicted in part (a) of Fig. 3, the lightweight ViT on the right is a scaled-down variant of the LPM on the left in terms of hidden dimensions and the number of layers, serving as the side network. The hidden dimensions of the side network are $1/k$ times that of the original, and its parameter weights are initialized randomly. Here, $k$ represents a reduction factor. The side network is shallower than the LPM in terms of the number of layers.
To utilize knowledge from the LPM, a novel fusion module $\mathcal{F}$ is introduced to fuse the intermediate features of the LPM and those of the side network at each layer, which will be elaborated upon in the subsequent section. Specifically, given the LPM $M$ with $L_{M}$ layers and the side network $S$ with $L_{S}$ layers, respectively parameterized by $\theta_{M}$ and $\theta_{S}$, the intermediate activation $\bm{z}_{S}^{l}$ of layer $l$ in the side network is obtained as follows:
$\bm{z}_{M}^{l^{\prime}}=\theta_{M}^{l^{\prime}}(\bm{z}_{M}^{l^{\prime}-1}),$ (4)
$\bm{z}_{S}^{l}=\theta_{S}^{l}(\mathcal{F}(\bm{z}_{S}^{l-1},\bm{z}_{M}^{l^{\prime}})),$ (5)
where $l^{\prime}=L_{M}-L_{S}+l$ and $l\in[1,L_{S}]$, and $\bm{z}_{S}^{L_{S}}$ is considered as the final output of the entire framework.
Adoption of the side network not only reduces the number of trainable parameters through its lightweight design but also mitigates memory expenses during the training phase. As shown in Fig. 4, given the separation between the side network and the pre-trained model, the forward pass of the frozen pre-trained model does not involve any learnable parameter. Consequently, updating the side network prevents the need for memory-intensive back-propagation inside the heavy LPM, resulting in notable memory savings during training.
Figure 4: There are no learnable parameters inside the LPM, eliminating the need for gradient computation inside the heavy LPM.
III-C Asymmetric Input
Fine-grained information holds significant importance in medical image analysis. To extract such detailed information, one straightforward approach involves increasing the resolution of the input image. However, due to the mechanics of input patchification and the self-attention layer in the ViT architecture, increasing the input resolution leads to a dramatic rise in memory consumption and training cost. For instance, in a ViT with a patch size of 16, increasing the input resolution from 224 to 512 results in an expansion of the input sequence length from 196 to 1024. Furthermore, the dimensions of the self-attention maps in the transformer block escalate from $196^{2}$ to $1024^{2}$. This considerable training overhead presents challenges for fine-tuning on high-resolution datasets.
Although high-resolution inputs make the training cost unfeasible, employing a pre-trained model for inference with high-resolution images remains practical. We here propose an asymmetric input strategy within the FPT+ framework. Specifically, the input images $I_{h}$ for the frozen pre-trained model are of high resolution, whereas the images fed to the learnable side network are low-resolution images $I_{l}$ that get down-sampled from the high-resolution ones. Therefore, the Equation 5 becomes
$\bm{z}_{S,I_{l}}^{l}=\theta_{S}^{l}(\mathcal{F}(\bm{z}_{S,I_{l}}^{l-1},\bm{z}_{M,I_{h}}^{l^{\prime}})).$ (6)
III-D Fine-grained Prompts and Fusion Modules
In this section, we introduce our proposed fusion modules $\mathcal{F}$ and fine-grained prompts. To fuse features from the LPM and those from the side network, a simple approach is to use an MLP layer to process and align the hidden dimensions of those features, and then concatenate them into the intermediate sequence of the side network. However, this approach will dramatically increase the length of the token sequence of the side network, significantly increasing memory consumption. Furthermore, popular LMPs are generally pre-trained on natural images, and since the pre-trained model remains frozen during training, integrating this out-of-domain knowledge might not seamlessly adapt to new medical downstream tasks. Moreover, directly fusing features across diverse scales and domains poses challenges for the side network’s ability to effectively learn from its input images.
To tackle these challenges, as illustrated in part (b) of Fig. 3, we introduce a small set of learnable embeddings $\bm{p}$ named fine-grained prompts and utilize a cross-attention mechanism [6, 46] to summarize pre-trained knowledge from the LPM. Unlike prompt tuning [31, 27] which directly uses prompts as part of the input sequence, our proposed fine-grained prompts serve as a bridge linking the frozen LPM and the side network. Specifically, in the context of cross-attention, we reuse the key $K_{M}$ and the value $V_{M}$ from the self-attention layer of the LPM, while the fine-grained prompts serve as the query. The cross-attention module allows the prompts to retrieve tokens of interest and summarize fine-grained information.
Subsequently, these fine-grained prompts are concatenated with the intermediate sequence of the side network, interacting with other tokens to share the fused pre-trained knowledge. The processing of the fusion module goes as follows:
$\mathcal{F}(\bm{z}_{S},\bm{z}_{M})$ $=[\bm{z}_{S},f_{out}(\text{CA}(\bm{p},\bm{z}_{M}))+\bm{p}],$ (7)
$\text{CA}(\bm{p},\bm{z}_{M})$ $=\text{Attn}(Q,K,V)$
$=\text{Attn}(\text{LN}(\bm{p}),K_{M},V_{M}),$ (8)
where notations for the input source and the layer index are omitted for simplification. The $\text{CA}(\cdot)$ and $\text{LN}(\cdot)$ denote the cross-attention module and the layer normalization operation, respectively. The [,][$\cdot$,$\cdot$][ ⋅ , ⋅ ] denotes the concatenation operation, and the out-projector $f_{out}$ is a linear layer to align the hidden dimension of the fine-grained prompts to that of the side network. The out-projector is shared across all fusion modules to further increase efficiency. Notably, the number of the fine-grained prompts is considerably smaller than that of the input sequence, leading to reduced training cost. They are removed after each layer’s forward processing in the side network to maintain consistency of the sequence length across layers.
III-E Important Tokens Selection
Despite no need for back-propagation from the frozen LPM and the great memory savings provided by the proposed fusion modules, integrating lengthy intermediate features produced by high-resolution inputs from the LPM still introduces a significant memory load. Inspired by the prior knowledge that medical images of the same modality often exhibit similar anatomical characteristics and that the diagnostic clues generally occupy only a small proportion of the entire image, we identify tokens containing diagnostic information within the input sequence of the pre-trained model as the main contributors for downstream tasks.
In such context, we present a straightforward yet effective approach to mitigate the training expenses: we select important tokens, and only features associated with those important tokens are passed to the fusion modules. Specifically, we assess the tokens’ importance based on the self-attention maps [15] in the LPM. The importance score of a token is defined as the average attention score of other tokens upon it. Subsequently, we select the top $m\%$ tokens with the highest scores as the important tokens, which are delivered to the side network through the fusion modules. This strategy alleviates the overhead introduced by the lengthy tokens from the high-resolution inputs, while preserving the discriminative features essential for downstream tasks.
III-F Fine-grained Features Preloading
To further enhance the training efficiency, we choose not to employ any data augmentation for the high-resolution input of the frozen LPM, ensuring that features associated with an image of interest remain constant throughout training. Consequently, it becomes feasible to pre-store these features prior to training, resulting in notable reductions in the training cost. Leveraging the benefits of important tokens selection, the demand for storage and the overhead linked to data loading are also considerably reduced, thereby collectively accelerating the training process. Although the high-resolution input of the LPM remains fixed, we apply data augmentation to the low-resolution input of the side network to maintain the diversity of the training samples. Note that the preloading technique is exclusive to FPT+, because the asymmetric input design allows for the application of preloading on the high-resolution input while avoiding overfitting and performance degradation through data augmentation on the low-resolution input.
TABLE I: Comparison results with state-of-the-art PETL methods across all evaluation datasets. ’Params’ denotes the ratio of learnable parameters to the total number of parameters. ’Mem.’ denotes the memory usage (MB). ’Avg.’, ’PPE’, and ’PME’ denote the average performance across all datasets, performance-parameter-efficiency, and performance-memory-efficiency metrics. The best and second-best results are indicated in bold and underlined formatting, respectively.
Computing cost Fundus image Dermoscopic image Mammography Chest X-ray
Method Params. $\downarrow$ Mem. $\downarrow$ DDR Messidor-2 ISIC2018 ISIC2016 DDSM CMMD COVID CHNCXR Avg. $\uparrow$ PPE $\uparrow$ PME $\uparrow$
Supervised (ImageNet)
Full fine-tuning 100 23,128 89.79 ± 0.76 86.87 ± 0.53 96.65 ± 0.29 83.22 ± 1.02 92.49 ± 0.34 66.26 ± 0.69 99.85 ± 0.05 95.40 ± 1.05 88.82 65.73 65.73
Linear probing 0.01 3,416 80.17 ± 0.85 79.73 ± 0.84 93.37 ± 0.31 81.27 ± 0.70 80.89 ± 0.39 60.23 ± 0.62 99.21 ± 0.07 91.34 ± 0.20 83.28 83.28 78.44
Prompt-tuning [27] 0.17 20,582 81.07 ± 0.59 80.02 ± 2.34 94.20 ± 0.20 81.66 ± 1.03 82.67 ± 0.34 59.62 ± 0.83 99.27 ± 0.03 90.95 ± 0.82 83.68 83.62 63.47
Attention-tuning [45] 33.8 20,792 81.99 ± 1.85 81.87 ± 1.42 94.40 ± 0.40 80.11 ± 1.92 80.67 ± 2.33 60.03 ± 1.03 99.58 ± 0.19 92.07 ± 2.48 83.68 73.74 63.34
Adapter [7] 2.03 19,360 85.84 ± 2.14 80.77 ± 7.48 95.96 ± 0.13 79.22 ± 2.26 80.76 ± 3.38 61.39 ± 1.40 99.17 ± 0.48 90.02 ± 0.91 84.14 83.41 64.61
BitFit [2] 0.12 20,382 84.31 ± 0.50 83.81 ± 1.11 94.84 ± 0.15 82.35 ± 0.62 84.77 ± 0.54 63.84 ± 1.80 99.81 ± 0.03 91.62 ± 0.95 85.67 85.63 65.11
LoRA [22] 0.68 20,970 85.84 ± 2.42 86.08 ± 0.95 95.02 ± 0.22 84.22 ± 1.98 82.26 ± 4.04 62.26 ± 1.66 99.69 ± 0.05 92.13 ± 1.12 85.94 85.69 64.93
AdaLoRA [51] 0.52 20,960 85.21 ± 0.34 86.07 ± 0.74 95.85 ± 0.24 84.23 ± 0.65 86.91 ± 0.30 62.45 ± 1.46 99.75 ± 0.04 92.25 ± 0.70 86.59 86.40 65.43
FPT [23] 1.81 1,824 84.13 ± 1.06 84.95 ± 2.01 93.88 ± 0.60 80.68 ± 1.13 90.52 ± 0.59 65.35 ± 0.94 99.70 ± 0.30 91.97 ± 1.36 86.40 85.73 83.60
FPT+ (Ours) 1.03 736 85.53 ± 1.58 86.31 ± 0.66 94.27 ± 0.32 82.91 ± 1.25 90.64 ± 0.71 66.33 ± 0.43 99.71 ± 0.18 91.30 ± 0.94 87.12 86.73 85.94
DINO (ImageNet)
Full fine-tuning 100 23,128 87.33 ± 0.77 84.06 ± 1.29 96.44 ± 0.26 84.46 ± 1.55 91.19 ± 0.32 64.70 ± 0.32 99.92 ± 0.03 95.31 ± 1.19 87.93 65.07 65.07
Linear probing 0.01 3,416 80.77 ± 0.67 81.71 ± 0.69 93.44 ± 0.16 83.96 ± 0.93 81.06 ± 0.46 63.64 ± 1.08 99.40 ± 0.01 91.84 ± 0.18 84.48 84.48 79.57
Prompt-tuning [27] 0.17 20,582 80.48 ± 0.34 80.71 ± 1.85 93.76 ± 0.11 83.41 ± 1.40 81.86 ± 0.34 63.30 ± 1.25 99.44 ± 0.14 91.97 ± 0.33 84.37 84.31 63.99
Attention-tuning [45] 33.8 20,792 82.38 ± 0.93 77.11 ± 3.29 92.60 ± 0.74 76.32 ± 4.05 75.58 ± 8.26 51.31 ± 2.63 99.57 ± 0.14 95.03 ± 0.30 81.24 71.59 61.49
Adapter [7] 2.03 19,360 87.95 ± 0.54 84.99 ± 1.89 96.02 ± 0.13 82.30 ± 2.26 77.61 ± 8.86 57.56 ± 1.48 99.79 ± 0.12 89.73 ± 4.55 84.49 83.76 64.88
BitFit [2] 0.12 20,382 86.35 ± 0.42 84.68 ± 0.42 95.11 ± 0.15 84.60 ± 1.13 84.27 ± 0.34 66.03 ± 1.53 99.82 ± 0.04 93.43 ± 0.64 86.79 86.74 65.96
LoRA [22] 0.68 20,970 85.14 ± 3.86 84.35 ± 1.27 95.97 ± 0.40 81.06 ± 6.82 $7$ 1.92 60.59 ± 2.36 99.89 ± 0.03 95.02 ± 0.70 84.97 84.72 64.20
AdaLoRA [51] 0.52 20,960 86.39 ± 0.37 84.27 ± 0.66 95.79 ± 0.25 83.74 ± 0.88 88.16 ± 0.35 64.39 ± 0.14 99.64 ± 0.08 93.29 ± 0.79 86.96 86.76 65.71
FPT [23] 1.81 1, 824 83.79 ± 0.76 84.69 ± 1.06 94.27 ± 0.18 83.05 ± 0.85 90.08 ± 0.68 66.23 ± 0.46 99.72 ± 0.05 91.55 ± 0.81 86.67 86.00 83.86
FPT+ (Ours) 1.03 736 86.52 ± 0.42 86.13 ± 0.85 93.82 ± 0.40 84.14 ± 1.65 90.50 ± 0.37 65.75 ± 0.44 99.73 ± 0.14 91.52 ± 0.72 87.26 86.87 86.08
IV Experiments
IV-A Datasets
We evaluate FPT+ using eight medical image datasets for classification tasks across four modalities, including fundus images (DDR [33] and Messidor-2 [11]), dermoscopic images (ISIC 2016 [18] and ISIC 2018 [9]), mammography images (DDSM [30] and CMMD [10]), and chest X-ray images (COVID [42] and CHNCXR [26]). The dataset sizes range from 662 to 12,522 samples, with classification categories varying from 2 to 7. When official splits for training, validation, and testing are available, they are adopted. Otherwise, a random partition of 70%/10%/20% is employed for training, validation, and testing, respectively.
IV-B Training and Evaluation Setup
IV-B1 Experiment Setup
In this study, we employ two popular variants of ViT [15], specifically ViT-B(ase) and ViT-L(arge), as our large-scale pre-trained models, with 86 million and 304 million parameters, respectively. To evaluate the generalizability of different PETL methods, we experiment with two types of pre-trained weights: one obtained through supervised learning and the other via self-supervised learning using DINO [5]. These models, which are pre-trained on the ImageNet-21K dataset [40], are widely recognized for their robust generalizability in natural image tasks. The patch size of ViTs used in this work is 16. All methods are trained for 20 epochs with a mini-batch size of 16 on an NVIDIA RTX A6000 GPU (48GB) using PyTorch of version 2.0.0 [36]. We adopt a cosine annealing schedule for the learning rate during training. ViTs are trained using the AdamW optimizer with cross-entropy as the loss function for all datasets. Each experiment is repeated 5 times, and we report the mean ± standard deviation values.
If not specified, all compared methods run at a resolution of $512\times 512$. In the context of FPT+, the high resolution for the LPM remains $512\times 512$, while the low resolution for the side network is set to $128\times 128$. The number of fine-grained prompts is set to 16, and the reduction factor $k$ is set to 8. For important tokens selection, the top 20% of the important tokens are retained for utilization within the fusion modules. The number of layers in the side network $L_{S}$ is set to 6, which is half the number of layers in the LPM.
For compared methods, LoRA [22] and AdaLoRA [51] use a scale factor ($\alpha$) of 16, with LoRA layers applied to the query and value weights in the ViT. Prompt tuning is reproduced following the VPT-deep design [27], with 16 prompts. To ensure fair comparisons across methods, grid searching is utilized to determine the optimal set of hyperparameters, including the learning rate and the weight decay, based on performance on the validation set. All methods and the side network of FPT+ are trained using the same data augmentations, including cropping, horizontal flipping, and color distortion.
IV-B2 Evaluation Metrics
We use the Area Under the Receiver Operating Characteristic Curve (AUC) to evaluate the classification performance on each dataset. To evaluate the performance and efficiency balance, we adopt the metrics proposed in our previous work [23], which extend from the performance-efficiency metric [32, 20]. Specifically, two metrics, namely performance-parameter-efficiency (PPE) and performance-memory-efficiency (PME), are utilized to assess the impact of the number of learnable parameters and GPU memory consumption on efficiency during transfer learning, respectively. They are defined as
PPE $=\text{score}\times e^{-\log_{10}(r+1)},$ (9)
PME $=\text{score}\times e^{-\log_{10}(m+1)},$ (10)
where score is the average performance across all datasets, $r$ is the ratio of learnable parameters to all parameters, and $m$ is the ratio of GPU memory requirement of a method of interest to that of fine-tuning the entire LPM.
TABLE II: Training efficiency comparisons across different PETL methods on the Messidor2 dataset.
FLOPs (GB) $\downarrow$
Total Training
time (s)
$\downarrow$
Throughput
(img/s)
$\uparrow$
Peak Power
Usage (W)
$\downarrow$
Avg. Power
Usage (W)
$\downarrow$
Full FT 87.84 1135.91 18.48 295.62 280.92
Linear 87.84 416.72 51.06 286.61 278.59
Prompt-tuning [27] 88.30 946.55 22.19 296.69 286.52
Attention-tuning [45] 87.84 1005.62 20.89 291.76 281.79
Adapter [7] 89.68 954.51 21.93 297.30 288.30
BitFit [2] 87.84 966.69 21.75 291.40 279.66
LoRA [22] 107.89 1014.68 20.62 295.85 288.47
AdaLoRA [51] 107.28 1027.17 20.44 292.00 280.56
Preloading 87.84 31.76 32.81 221.812 217.74
FPT 0.28 154.23 171.83 166.54 183.10
FPT+ 0.06 90.61 357.37 114.38 107.60
TABLE III: Comparison results with state-of-the-art PETL methods using a ViT-L backbone with supervised pre-trained weights, evaluated at different input resolutions. The ’/’ symbol indicates out-of-memory conditions on a 48GB GPU. The average AUC is computed across all eight classification tasks.
Memory Usage $\downarrow$ Average AUC $\uparrow$
Method 128 256 384 512 640 128 256 384 512 640
Full FT 7,816 12,168 26,470 / / 83.94 86.96 87.23 / /
Linear probing 1,640 1,884 2,778 5,066 9,716 77.46 82.99 84.38 84.62 84.77
Prompt-tuning [27] 3,128 7,602 20,392 / / 79.82 83.90 84.95 / /
Attention-tuning [45] 4,340 8,710 21,516 / / 80.93 82.21 83.89 / /
Adapter [7] 3,248 7,608 20,394 / / 79.30 81.70 82.88 / /
BitFit [2] 3,058 7,186 20,074 / / 82.85 85.46 86.38 / /
LoRA [22] 3,152 7,810 21,190 / / 80.13 82.96 83.95 / /
AdaLoRA [51] 3,142 7,806 21,056 / / 82.17 85.52 86.46 / /
FPT [23] 2,546 2,724 3,152 3,730 4,456 80.72 84.34 85.75 86.85 87.28
FPT+ (Ours) 570 630 718 862 1,050 81.74 85.60 86.88 87.87 88.14
IV-C Comparisons with State-of-the-art
We compare FPT+ against full fine-tuning, linear probing, and SOTA PETL approaches, including our previous work FPT. In full fine-tuning, all parameters are learnable when training on downstream tasks. Although full fine-tuning often represents the performance’s upper bound for fine-tuning, the relative scarcity of training samples in medical image datasets may lead to overfitting and consequent performance degradation. Linear probing emerges as an efficient strategy to adapt pre-trained models to downstream tasks by solely training a new task-specific head on top of the pre-trained model. Additionally, we compare FPT+ with popular SOTA PETL methods applicable to vision tasks.
The performance comparisons are tabulated in Table I. Full fine-tuning presents strong performance across all datasets but comes with the highest computational cost, requiring a 100% parameter ratio and significant memory usage (23,128 MB), making it impractical for resource-constrained settings. Linear probing, on the other hand, shows low parameter usage (0.01%) and memory consumption (3,416 MB). However, this method generally underperforms. Other PETL methods are more balanced between parameter efficiency and performance. For example, LoRA and AdaLoRA, with 0.68% and 0.52% parameter ratios respectively, show competitive performance across various datasets.
Although these aforementioned methods achieve reductions in the number of learnable parameters, their memory consumption remains high. For instance, Adapter has the lowest memory consumption but still requires 19,360 MB to fine-tune the LPM. In contrast, our proposed method, FPT+, outperforms other PETL methods in both performance and efficiency, achieving the best performance among PETL methods for both pre-trained weights and the best performance-efficiency balance (PPE and PME) among all methods. Specifically, FPT+ exhibits remarkable parameter efficiency with only 1.03% of learnable parameters and significantly lower memory consumption (736 MB, 3.18% compared to full fine-tuning), without compromising performance. Our previous work, FPT, already shows significant improvements in parameter efficiency and performance. On top of FPT, FPT+ further enhances these metrics, achieving higher performance and reduced memory consumption (736 MB vs. 1,824 MB).
The comparison results emphasize the effectiveness of FPT+ in balancing performance, parameter efficiency, and memory consumption. Its design ensures efficient high-resolution medical image classification, making it a promising solution for advancing medical diagnostics. The generalizability and robustness of FPT+ are further highlighted by its effectiveness with both supervised and DINO pre-trained weights, making it adaptable to different pre-training paradigms and practical for a wide range of medical image classification tasks.
IV-D Training Efficiency
To comprehensively evaluate the efficiency of PETL methods, we further assess several metrics related to training cost, including computational cost (FLOPs), total training time, training time per epoch, throughput, and power usage (both peak and average). These training efficiency metrics are evaluated using the Messidor2 dataset, containing 1,748 samples with a resolution of 512 × 512, and employing ViT-B as the backbone. Table II provides a detailed comparison of these metrics across various methods.
As shown in Table II, FPT+ demonstrates remarkable training efficiency with computational cost of only 0.06 GB FLOPs, being significantly lower than full fine-tuning (87.84 GB) and PETL methods like LoRA (107.89 GB). FPT+ also excels in terms of throughput, processing 357.37 images per second, which is a significant improvement over other PETL methods like LoRA (20.62 images/second) and Prompt-tuning (22.19 images/second). Compared to FPT, FPT+ doubles the throughput, showing its enhanced efficiency. With these reductions in the computational cost, FPT+ achieves a total training time of 90.61 seconds, which is more than 12 times faster than full fine-tuning (1135.91 seconds). FPT+ is also energy-efficient, achieving an average power usage of 107.60 W and a peak power usage of 114.38 W, staying far below the average of 280 W and the peak usage of over 290 W for other PETL methods on GPUs with a 300 W power limit.
Additionally, FPT+ incorporates a preloading step, which extracts and pre-stores features from the frozen LPM before training. This step, taking only 31.76 seconds, adds negligible overhead compared to the substantial training costs of other fine-tuning methods.
These results establish FPT+ as a highly efficient PETL method in terms of computational, temporal, and energy resources, making it a practical and effective solution for high-resolution medical image classification tasks, especially in resource-constrained settings.
IV-E Evaluation with A Larger Pre-trained Model
To further investigate the generalizability of FPT+ and highlight its significant advantage in memory saving, we evaluate its performance and efficiency by comparing it with SOTA methods using a ViT-L backbone with supervised pre-trained weights across different input resolutions. The superiority of FPT+ becomes even more evident when using a larger model. As shown in Table III, most PETL methods encounter out-of-memory issues on a 48GB GPU when the input resolution exceeds 384 $\times$ 384, limiting their capability to improve performance through higher resolutions. In contrast, FPT, FPT+, and linear probing, with better memory efficiency, can run at higher resolutions and exhibit greater performance potential.
Full fine-tuning achieves its best performance with an average AUC of 87.23, at its highest feasible resolution of 384 $\times$ 384. Similarly, the PETL method AdaLoRA achieves its best performance with an average AUC of 86.48, also at the resolution of 384 $\times$ 384. These methods are constrained by memory limitations, which largely restrict their feasibility to operate at higher resolutions. Although linear probing can run at a resolution of 640 $\times$ 640, its performance remains relatively low, with an average AUC of 84.77 at the resolution of 640 $\times$ 640. In contrast, FPT+ not only achieves the best performance among all PETL methods at the resolution of 384 $\times$ 384 with an average AUC of 86.88 but also runs efficiently at the higher resolution of 640 $\times$ 640, achieving the highest overall performance with an average AUC of 88.14. These results demonstrate the superior scalability of FPT+ when handling higher resolutions.
One of the key advantages of FPT+ is its exceptional memory efficiency. At a resolution of 640 $\times$ 640, FPT+ has a memory usage of 1,050 MB, which is even lower than the memory usage of other methods at a resolution of 128 $\times$ 128. Such memory efficiency enhances the feasibility of deploying FPT+ and enables the exploration of higher input resolutions to potentially attain even better performance.
Figure 5: Impact of the number of layers on FPT+’s performance and efficiency. The star symbol indicates the best performance.
Figure 6: Average attention score on the fine-grained prompts at each layer.
IV-F Contribution of Each Layer
In this section, we analyze the behavior of our previous work, FPT, in terms of fine-grained prompts and show how FPT+ empirically simplifies FPT’s architecture. As described in FPT, similar to FPT+, the pre-trained knowledge is conveyed to the side network via fine-grained prompts and fusion modules. The fine-grained prompts are concatenated with the intermediate sequence of the side network to join the forward propagation. The side network integrates the pre-trained knowledge through self-attention layers, where other tokens obtain information from the fine-grained prompts by paying attention to these prompts. However, unlike FPT+, which has a side network of $L_{S}$ layers, in FPT, each layer in the LPM has a corresponding learnable layer and a corresponding fusion module, resulting in a side netowrk of $L_{M}$ layers, which induces non-trivial computational overhead.
Since higher layers of a ViT-based LPM are more likely to contain high-level information for image understanding [39], we conjecture that not all layers of the side network are necessary for high performance. To investigate the contribution of each layer to downstream tasks, we visualize the average attention score that other tokens pay to the fine-grained prompts at each layer of FPT. As shown in Fig. 6, the fusion modules pay more attention to the fine-grained prompts in higher layers; the average attention score of layers 7-12 is 0.315, while that of layers 1-6 is 0.180. This suggests that leveraging information from higher layers is more critical for downstream tasks, allowing for a more efficient design of the side network by focusing on key layers.
To determine the number of LPM layers necessary to provide sufficient information for downstream tasks, we sequentially add layers to the side network starting from the highest to the lowest, including the corresponding fusion modules, and assess their impact. As presented in the right part of Fig. 5, the average AUC increases with the number of layers and saturates at 6 layers, demonstrating that the pre-trained information from the highest 6 layers is sufficient for downstream tasks. Although FPT+ with 9 layers achieves the best average AUC, increasing the number of layers also increases the number of learnable parameters and memory consumption. Furthermore, FPT+ with 6 layers achieves very close AUC compared to 9 layers (87.12 vs 87.32), and the best performance-efficiency balance in terms of PPE and PME. Therefore, we set the default number of layers $L_{S}$ for the side network to 6.
IV-G Ablation Study
In this section, we systematically investigate the impact of each component in FPT+. Unless otherwise specified, the presented performance metric is the average AUC across eight datasets using the pre-trained weights in a supervised manner, and the configuration of FPT+ remains the same as that described in Section IV-B1.
IV-G1 Impact of Components
To assess the contributions of individual components within our proposed FPT+, we conduct an ablation study by adding components sequentially and measuring their impact on AUC and GPU memory usage. As shown in Fig. 7, each additional component incrementally enhances the model’s performance, with the average AUC improving from 74.74 to 87.12. The most substantial performance gain is attributed to the introduction of pre-trained knowledge from the LPM. However, simply using MLP for transferring pre-trained knowledge significantly increases the memory consumption from 8,214 MB to 35,682 MB. Replacing MLP with fusion modules notably enhances memory efficiency by reducing the memory consumption from 35,682 MB to 13,328 MB, followed by consistent improvements with the incorporation of asymmetric inputs, important tokens selection, and features preloading down to 736 MB. By combining these proposed components, FPT+ achieves the optimal balance of high performance and low memory consumption.
Figure 7: Impact of each component on FPT+’s performance and efficiency. The average AUC and GPU memory usage are shown for each configuration, starting with the side network alone and sequentially adding components.
IV-G2 Input Resolutions
To reduce memory consumption, we propose asymmetric inputs for the LPM and the side network, where the input for the side network is a down-scaled version of that for the LPM. In this section, we investigate the impact of the input resolution for the side network. We train FPT+ using different input resolutions for the side network while maintaining the input resolution for the LPM at $512\times 512$. As shown in Table IV, the results indicate that a resolution of $128\times 128$ yields optimal performance while maintaining memory efficiency. We conjecture that the poor performance observed with a higher down-scaled resolution could potentially be attributed to the lengthy input sequence in the side network, which distracts attention from the fine-grained prompts. Notably, even using an input resolution of $64\times 64$ for the side network achieves better performance compared to other PETL methods, demonstrating that the proposed fine-grained prompts and fusion modules can effectively convey fine-grained information for diagnostics in the side network.
TABLE IV: FPT+’s performance and efficiency with different input resolutions for the side network.
Input Resolution Avg. AUC $\uparrow$ Mem. $\downarrow$ PME $\uparrow$
64 86.78 694 85.67
128 87.12 736 85.94
256 86.32 1,342 84.23
512 84.78 8,704 73.80
IV-G3 Number of Fine-grained Prompts
The number of fine-grained prompts can be significantly fewer than the total tokens from the input images presented in both the pre-trained model and the side network, thereby reducing the memory consumption. We here explore the influence of the number of fine-grained prompts. As depicted in Table V, with the increase in the number of prompts, the performance improves and the best performance is achieved at 16 prompts, which number is sufficient to summarize fine-grained information and convey it to the side network. We also notice that even with only 4 prompts, the performance is comparable to other PETL methods, demonstrating the effectiveness of FPT+’s fusion modules in obtaining pre-trained knowledge from the LPM.
TABLE V: FPT+’s performance and efficiency with different numbers of prompts for the side network.
# Prompt Avg. AUC $\uparrow$ Param. $\downarrow$ Mem. $\downarrow$ PPE $\uparrow$ PME $\uparrow$
4 85.94 0.96 720 85.58 84.80
8 86.39 0.99 722 86.02 85.24
16 87.12 1.03 736 86.73 85.94
32 86.54 1.11 812 86.13 85.25
64 86.66 1.28 972 86.18 85.12
128 86.49 1.61 1,388 85.89 84.33
IV-G4 Reduction Factor
The reduction factor $k$ in FPT+ determines the hidden dimension of the side network. As shown in Table VI, we assess the impact of different reduction factors on FPT+’s performance and efficiency. A reduction factor of 8 achieves the highest average AUC (87.12) with minimal memory usage (736 MB) and only 1.03% of learnable parameters. Increasing the reduction factor to 16 slightly reduces memory usage but results in a drop in AUC (85.92), while lower factors (e.g., 2 and 4) increase resource consumption without significant gains in performance. We observe that a larger hidden dimension in the side network (lower reduction factor) does not improve performance as it increases the training complexity of the randomly initialized side network. A reduction factor of 8 achieves the optimal balance for high-resolution medical image classification tasks.
TABLE VI: FPT+’s performance and efficiency with different reduction factors.
Reduction Factor Avg. AUC $\uparrow$ Param. $\downarrow$ Mem. $\downarrow$ PPE $\downarrow$ PME $\downarrow$
2 86.32 11.55 1,170 82.32 84.49
4 86.74 3.40 894 85.49 85.32
8 87.12 1.03 736 86.73 85.94
16 85.92 0.37 724 85.78 84.78
IV-G5 Ratio of Important Tokens
To mitigate the training cost, we exclusively integrate features derived from important tokens of the pre-trained model into the side network. In this section, we assess how different ratios of these important tokens affect the process. We provide details on the average performance and memory usage of FPT+ across various token ratios. We notice that the optimal ratio of important tokens varies across different datasets. For instance, although both DDR and Messidor-2 are fundus image datasets, they show peak performance at different important token ratios. Although a higher ratio of important tokens can provide more information for fusion modules, it carries a risk of misleading the fusion modules and potentially increasing the difficulty of learning. Most datasets prefer a token ratio lower than 0.5, except for Messidor-2 and CHNCXR which achieve the best performance at 0.9 and 0.7, respectively. We conjecture that the optimal token ratio depends on the scale and complexity of the dataset of interest.
In terms of memory efficiency, as the ratio of important tokens increases, the memory consumption also rises dramatically, resulting in significant drops in PME. Overall, a token ratio of 0.2 emerges as the optimal choice. At this ratio, FPT+ achieves the best average AUC with low memory consumption, ensuring an efficient balance. The PME metric also indicates that a 0.2 ratio provides the best balance between performance and memory efficiency.
Figure 8: Impact of important tokens’ ratio on FPT+’s performance and efficiency. The star symbol indicates the best performance. Increasing the ratio does not affect the number of learnable parameters, resulting in a constant PPE (represented as a horizontal line).
IV-H Visualization
To verify the important tokens selection and the mechanism of the fusion modules, we visualize the selected tokens of the top layer and the attention scores on them in the corresponding fusion module across four modalities. As shown in Fig. 9, the selected important tokens are highlighted in green boxes. It can be observed that the selected tokens cover most discriminative regions, such as lesions and the optic disc in fundus images. Please note that ViT employs a multi-head mechanism, and the visualization is only for one of the multiple heads. This multi-head mechanism ensures that the important tokens selection can convey sufficient diagnostic clues to the side network. Additionally, we visualize the attention that the fine-grained prompts pay to the selected tokens. It can be observed that the fine-grained prompts further refine the selection, reducing redundancy and summarizing them into a small set of prompts for the side network. This refinement process ensures that only the most relevant information is conveyed, enhancing the efficiency and effectiveness of the transfer learning process.
Figure 9: Visualization of the selected important tokens and attention scores from the fine-grained prompts.
V Discussion
With the success of large-scale pre-trained models, transfer learning has become a key technique to apply pre-trained knowledge in real-world applications. PETL has been proposed to efficiently adapt LPMs to new downstream tasks. However, existing PETL methods mainly focus on reducing the number of learnable parameters, and the improvements in memory efficiency are not as effective as those in parameter efficiency. Medical image analysis is a domain that often requires high-resolution inputs to ensure that tiny diagnostic clues are clearly exposed. However, increasing the input resolution dramatically increases memory consumption during training. As shown in Fig. 2, even the most memory-efficient PETL method, linear probing, is not feasible when the input resolution is $1280\times 1280$ on a GPU with a 48GB memory size. Our proposed FPT+ significantly improves memory efficiency, making training at $2048\times 2048$ possible on a single GPU, with the potential for even higher resolutions based on the trend shown in the figure.
Although FPT+ makes the training of super high-resolution images feasible on one GPU, we notice that the performance saturates or even slightly drops when the input resolution exceeds $1024\times 1024$. There may be two plausible reasons: (1) The LPM used is pre-trained at a standard resolution of $224\times 224$, and its feature extraction ability degrades when the resolution is significantly different from the pre-trained resolution; (2) The datasets we use may have already been resized and compressed by their owners, allowing users to download them from the website. Hence, increasing the input size to go beyond the original image size shall not provide any additional information.
High-resolution image classification is not the only field that benefits from this memory efficiency. FPT+ can also be applied to ViT variants with smaller patch sizes. ViTs with smaller patch sizes generally have more powerful abilities to extract fine-grained semantics. However, smaller patch sizes result in longer input sequences for the same input resolution. FPT+ may be a good solution to attain the best balance. The memory efficiency of FPT+ also allows for larger batch sizes during training, which can accelerate the training process and potentially improve the stability and performance of transfer learning. Additionally, there are settings in natural image analysis that also require high resolution, such as satellite imagery [1]. The generalizability of FPT+ to natural images remains unexplored, which is one of our future works.
VI Conclusion
In this paper, we extend our previous work FPT [23] to FPT+. FPT+ significantly improves the memory efficiency over existing PETL methods, particularly in the high-resolution context commonly encountered in medical image analysis. To reduce memory consumption, we first enhance the design of side tuning with asymmetric input. To effectively adapt pre-trained knowledge from the LPM, we introduce fine-grained prompts and fusion modules to summarize and process intermediate features of the LPM and effectively convey them to the side network. To further reduce memory utilization, we propose the selection of important tokens and the preloading of pre-trained features. By integrating all these components, we present a parameter and memory-efficient transfer learning method for high-resolution medical image classification. The effectiveness of FPT+ is successfully established across eight medical image datasets of diverse sizes, modalities, and complexities. Experimental results show that FPT+ outperforms all other SOTA PETL methods while maintaining the best parameter and memory efficiency.